OpenStudy (anonymous):

More help? Type the equation of the given line in standard form. The line with m = 4 and intersecting 2x -4y = 8 at x = 6. Thanks!

OpenStudy (anonymous):

Step #1: convert "2x -4y = 8" to y=mx+b Step #2...

OpenStudy (anonymous):

So that would make it -4y=2x+8?

OpenStudy (anonymous):

Or do I have to put the m in there to and make it -4y=8x+8?

OpenStudy (anonymous):

you want y not -4y so you should simplify

OpenStudy (anonymous):

Move the 2x over to the other side (with subtraction), divide both sides by -4. Can you write this second step @cutie.patootie ? :-) You'll end up getting a y=mx+b form line that allows to to draw this: